# RAS: Retrieval-And-Structuring for Knowledge-Intensive LLM Generation

### v1

@@ Abstract
Retrieval-augmented language models often struggle with knowledge-intensive tasks due to inefficient retrieval, unstructured knowledge integration, and single-pass architectures. We present Retrieval-And-Structuring (RAS), a novel framework that dynamically constructs and reasons over query-specific knowledge graphs through iterative retrieval and structuring. RAS introduces four key technical innovations: (1) a theme-scoped retrieval mechanism that efficiently narrows the search space while maintaining retrieval quality, (2) an action planning module that determines knowledge needs and generates focused sub-queries, (3) a dynamic knowledge structuring approach that converts retrieved text into an evolving knowledge graph, and (4) a graph-augmented answering component that leverages the accumulated structured information. Our framework achieves state-of-the-art performance, surpassing leading baselines by 6.4% with open-source language models and 7.0% with proprietary models on seven knowledge-intensive generation datasets across all evaluation metrics. Detailed ablation studies verify the contribution of each technical component to the overall system performance.
@@ 1 Introduction
Complex reasoning tasks like scientific analysis or multi-hop question answering demand both comprehensive knowledge and structured logical thinking. While large language models (LLMs) have achieved remarkable performance across natural language processing tasks (Devlin et al., 2018; Brown et al., 2020; Chowdhery et al., 2023), they often struggle with knowledge-intensive reasoning due to the absence of precise, logically structured information (Rae et al., 2021; Ling et al., 2024). This limitation has prompted growing research into augmenting LLMs with structured knowledge to enhance their reasoning capabilities (Wang et al., 2021).
Retrieval-augmented generation (RAG) approaches have emerged as a promising direction by providing LLMs with additional context from retrieved passages (Guu et al., 2020; Lewis et al., 2020; Izacard & Grave, 2021; He et al., 2024). However, these methods frequently face hallucination challenges (Maynez et al., 2020; Zhang et al., 2023b), where generated content deviates from retrieved information. This issue stems primarily from the unstructured nature of retrieved passages, which lack explicit logical connections needed for precise reasoning. Recent efforts have explored integrating knowledge graphs (KGs) with LLMs (Sun et al., 2019; Yu et al., 2022; He et al., 2024; Edge et al., 2024), as KGs offer compact, structured representations of entities and relationships that enable more accurate and interpretable reasoning (Hogan et al., 2021; Jiang et al., 2024).
We present Retrieval-And-Structuring (RAS), a novel framework that dynamically constructs and reasons over query-specific knowledge graphs through iterative retrieval and structuring. RAS integrates theme-aware retrieval, dynamic KG construction, and self-refinement (Ji et al., 2023; Asai et al., 2023; Lyu et al., 2024b) to comprehensively prepare knowledge for challenging questions. The framework comprises four key components: (1) an action planner that strategically determines knowledge needs and generates targeted sub-queries, (2) a theme-scoped retrieval module that efficiently combines thematic and semantic relevance, (3) a knowledge structuring module that builds and expands query-specific KGs, and (4) a knowledge-augmented answerer that leverages the gathered structured information.
RAS addresses several critical limitations of existing approaches. First, unlike traditional RAG methods that rely solely on semantic similarity (Lewis et al., 2020; Guu et al., 2020), RAS employs theme-aware retrieval that significantly reduces the search space while preserving retrieval quality. Second, in contrast to static KG approaches (He et al., 2024; Edge et al., 2024), RAS dynamically constructs and evolves KGs based on the specific reasoning requirements of each query. Third, through its iterative planning mechanism, RAS actively identifies knowledge gaps and initiates focused retrievals, moving beyond the single-pass limitations of existing methods (Izacard & Grave, 2021).
@@
On seven knowledge-intensive tasks including open-domain QA, closed-set QA, and long-form generation, RAS outperforms state-of-the-art baselines by 6.4% with open-source LLMs and 7.0% with proprietary models. The main contributions of this work:
@@
We introduce RAS, a novel framework that dynamically constructs query-specific knowledge graphs through iterative retrieval and structuring for knowledge-intensive language generation.
We propose theme-scoped retrieval that combines thematic classification with dense retrieval to efficiently narrow the search space while maintaining retrieval quality.
We develop a dynamic knowledge structuring approach that iteratively builds and enriches knowledge graphs based on reasoning needs.
@@
We validate RAS’s effectiveness through comprehensive experiments and ablation studies across diverse knowledge-intensive tasks.
@@ 2 Related Work
Retrieval-Augmented Generation. Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020) enhances performance on knowledge-intensive tasks by incorporating retrieved passages into input of large language models (LLMs), improving factual accuracy and grounding. Traditional approaches retrieved a fixed number of passages once before generation (Shao et al., 2023; Es et al., 2024; Lyu et al., 2024a) or focused on adaptive retrieval (Jiang et al., 2023), where models dynamically query external knowledge when confidence is low, or analyze on the retrieval (Kim et al., 2024b). However, these methods often fail to perform complex or multi-hop reasoning process due to the linear and simple designs, where several recent methods have emerged to improve retrieval quality. Iterative retrieval-generation approaches (Shao et al., 2023; Guan et al., 2024) leverage historical context to enhance subsequent retrievals. Other work introduces targeted retrieval strategies through subqueries (Khattab et al., 2023) or follow-up questions (Yao et al., 2023; Press et al., 2022), progressively enriching the context for more factually grounded generation. SelfRAG (Asai et al., 2023) introduced self-reflective RAG, which allows LLMs to retrieve content on demand and critically evaluate retrievals, demonstrating both the importance of retrieval quality and the capacity of LLMs to seek self-directed information. RPG (Lyu et al., 2024b) further refined this approach by extracting fine-grained paragraphs from retrieved content to improve query relevance. However, existing methods face key limitations. They often either include irrelevant content in the retrieved context or risk discarding essential information through refinement. Additionally, dense retrieval approaches can be computationally expensive when deployed on large corpora, hampering real-time user interaction. To address these challenges, we propose converting retrieved text into knowledge-intensive graphs and maintaining an iteratively enriched, query-specific knowledge graph for knowledge serving. We also introduce theme-scoped retrieval for more focused and efficient context retrieval.
Graph as Context for LLMs. The integration of graphs into modern language models represents a promising research direction (Yasunaga et al., 2021, 2022; Yu, 2022; Ju et al., 2022; Zhu et al., 2024; Gutiérrez et al., 2024). Recent approaches leverage graph-based knowledge representation to enhance LLM performance. For instance, GraphToken (Perozzi et al., 2024) shows that LLM can understand a graph encoded as a token, after a simple p-tuning (Liu et al., 2021) while keeping the LLM’s parameters frozen. Based on the findings, G-Retriever (He et al., 2024) constructs subgraphs by retrieving relevant entities and relations from a global KG, using these as structured knowledge input for LLMs. GraphRAG (Edge et al., 2024) employs an alternative strategy, extracting a comprehensive global KG from the corpus, identifying meaningful clusters through graph community detection, and generating LLM-based summaries of these communities to create a refined retrieval corpus. These methods demonstrate how graphs, with their knowledge-intensive modality and inherent structural properties, can significantly improve LLMs’ contextual understanding. However, current approaches treat graphs as static data structures without refinement based on LLM requirements. Moreover, building global KGs for large-scale corpora remains prohibitively expensive and often impractical – Wikipedia alone contains over 30 million documents per dump. The key challenge lies in developing methods to dynamically construct and utilize graphs on-the-fly for knowledge-intensive generation. To address these limitations, we propose a retrieval-and-structuring paradigm that dynamically builds and maintains query-specific KGs on-the-fly. Our approach iteratively enriches these graphs through LLM’s self-evaluation of information completeness and targeted retrieval via focused subqueries.
3 RAS Framework
Knowledge-intensive language generation requires not just retrieving relevant information, but also structuring and reasoning over it effectively. We propose Retrieval-And-Structuring (RAS), a framework shown in Figure 1. In the following subsections, we detail the technical mechanisms underlying each stage. To facilitate understanding, we showcase a running example of RAS in Figure 26.
3.1 Action Planning
The Action Planning stage is the first critical step in the RAS framework, guiding the retrieval and reasoning process based on the input query and the current knowledge state. It involves two key components: the Action Planner , a language model that assesses the sufficiency of the current knowledge state and generates sub-queries for further retrieval if necessary, and a decision step that determines the next action based on the planner’s output.
@@ 3.1 Action Planning
(For initial iteration) Formally, given an input query , the Action Planner generates a plan according to:
where can take one of two forms:
[NO_RETRIEVAL]: If determines that can be answered directly without requiring any additional knowledge, the planning process terminates, and the framework proceeds directly to the final Answer Generation stage (§3.4).
@@
(For iteration 0) At each iteration , the Action Planner reassesses the cumulative retrieved knowledge and the chain of previous sub-queries in the context of the main query :
@@
where is a graph neural network for encoding the evolving KG ; is the subquery at iteration , and is the extracted graph information (triple list) from the retrieved context. The base model we use here is GraphLLM, an architecture derived from previous work (Perozzi et al., 2024; He et al., 2024). We discuss more details in §3.3.
The planner’s output at each iteration can be either:
[SUFFICIENT]: The accumulated knowledge is deemed sufficient to comprehensively address the main query . The iterative retrieval process terminates, and the framework proceeds to the Knowledge-Augmented Generation stage (§3.4).
@@
[SUBQ] : The planner generates a new subquery to guide the retrieval of additional relevant knowledge. The subquery is designed to fill specific gaps in the current knowledge state with respect to answering . The framework continues the iterative process and proceeds to the Theme-Scoped Retrieval stage (§3.2).
@@
The Action Planner serves as a key driver of the RAS framework’s iterative retrieval and refinement process. By dynamically assessing the adequacy of the retrieved knowledge and generating targeted sub-queries, it enables the efficient acquisition of query-relevant information.
3.2 Theme-Scoped Retrieval
This stage efficiently retrieves thematically and semantically relevant context for subquery through a two-step process: (1) theme scoping to identify a narrowed sub-corpus close to ’s thematic distribution, and (2) dense retrieval to find the most relevant passages within . This thematic pre-filtering before semantic matching aims to improve both retrieval efficiency and precision.
3.2.1 Theme Scoping
The Theme Scoping mechanism transforms sub-queries into thematic probability distributions and retrieves relevant document subsets through three key operations:
First, a theme classifier maps the subquery to a posteriori probability distribution (PPD) over thematic classes (Eq. 3). The classifier is trained on DBPedia-298 (Lehmann et al., 2015)111Document label space of DBPedia-298: https://mappings.dbpedia.org/server/ontology/classes/, a multilabel wiki text classification dataset, enabling fine-grained thematic distinctions.
To address the distribution shift between queries and documents, an MLP-based shifter transforms the query theme distribution into an expected document theme distribution (Eq. 4). This shifter is trained on query-document pairs in our constructed dataset HotpotQA-SUBQ (see Appendix B), learning to capture the relationship between query and document thematic patterns.
Finally, using the shifted distribution , we retrieve a theme-specific sub-corpus from the theme-indexed corpus (Eq. 5). The function selects the top- documents based on L2 similarity between and the pre-computed theme distributions in . This focused retrieval process effectively reduces the search space while maintaining thematic relevance for subsequent semantic matching operations.
3.2.2 Dense Retrieval
Within the theme-specific sub-corpus , a dense retriever is employed to identify the most semantically relevant passages for the subquery :
where selects the top- passages based on the cosine similarity between the query embedding and the passage embeddings precomputed in the dense index.
3.3 Knowledge Structuring and Enrichment
@@
This stage is designed to transform the retrieved text passages into a structured graph representation and iteratively expand the query-specific knowledge graph .
3.3.1 Text-to-Triple Conversion
To capture the essential facts from the retrieved passages , a text-to-triples model is employed. We train the model on the full set of WikiOFGraph dataset (Kim et al., 2024a), an LLM-curated high-quality text-triples dataset, enabling it to generate quality and information-rich triples in the format:
@@ 3.3.1 Text-to-Triple Conversion
where each triple represents a subject-predicate-object fact extracted from the text. This structured representation allows for efficient downstream reasoning and simple external knowledge graph integration.
3.3.2 Iterative Knowledge Enrichment
@@
The extracted triples are then converted into a graph structure , where and denote the sets of nodes and edges, respectively. Each node corresponds to a unique subject or object entity in , while each edge represents a predicate connecting two entities. To enrich the graph with semantic information, the attributes of nodes and edges are obtained through a text encoder (Sentence-BERT (Reimers, 2019) in our case):
@@
To progressively expand the knowledge base in response to the evolving sub-queries, the structured graph at each iteration is merged into a evolving knowledge graph specific to the main query :
After enriching with the new knowledge, the action planner (§3.1) reassesses the current knowledge to determine the next step. Based on and the chain of previous subqueries and their associated graph information, the planner decides whether to generate another focused subquery for additional retrieval (§3.2) or to proceed with answering (§3.4) if the accumulated knowledge is deemed sufficient.
3.4 Knowledge-Augmented Generation
The final stage generates an answer to the main query using the Answerer model conditioned on knowledge graph and subquery chain .
@@ 3.4 Knowledge-Augmented Generation
attends to relevant knowledge in and subquery chain to generate accurate, coherent answers.
Note: In this study, and are two roles played by the same model with different instructions (see §C.3).

### v2

@@ Abstract
Large language models (LLMs) have achieved impressive performance on knowledge-intensive tasks, yet they often struggle with multi-step reasoning due to the unstructured nature of retrieved context. While retrieval-augmented generation (RAG) methods provide external information, the lack of explicit organization among retrieved passages limits their effectiveness, leading to brittle reasoning pathways. Recent interpretability studies highlighting the importance of structured intermediate reasoning further align with this perspective. We propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs query-specific knowledge graphs through iterative retrieval and structured knowledge building. RAS interleaves targeted retrieval planning with incremental graph construction, enabling models to assemble and reason over evolving knowledge structures tailored to each query. On seven knowledge-intensive benchmarks, RAS consistently outperforms strong baselines, achieving up to 6.4% and 7.0% gains with open-source and proprietary LLMs, respectively. Our results demonstrate that dynamic, query-specific knowledge structuring offers a robust path to improving reasoning accuracy and robustness in language model generation. Our data and code can be found at https://github.com/pat-jj/RAS.
@@ 1 Introduction
Complex reasoning tasks such as scientific analysis or multi-hop question answering demand both comprehensive knowledge and structured logical thinking [1]. While large language models (LLMs) have achieved remarkable performance across a wide range of natural language processing tasks [2, 3, 4], they often struggle with knowledge-intensive reasoning due to the absence of precise, logically organized information [5, 6]. This limitation has motivated growing research into augmenting LLMs with structured knowledge to enhance their reasoning capabilities [7].
Retrieval-augmented generation (RAG) approaches have emerged as a promising direction by providing LLMs with additional context from retrieved passages [8, 9, 10, 11]. However, these methods frequently face hallucination challenges [12, 13], where generated content deviates from retrieved information. This issue stems primarily from the unstructured nature of retrieved passages, requiring the model to implicitly bridge gaps during generation. Recent research in model interpretability further highlights this issue [14]. Attribution graph analyses have revealed that LLMs internally attempt multi-hop reasoning by chaining facts across different parts of context, but failures in these reasoning pathways are correlated with hallucinations and factual inconsistencies. These insights suggest that explicitly providing structured intermediate knowledge could significantly enhance the reliability and transparency of LLM reasoning. Recent efforts have explored integrating knowledge graphs (KGs) with LLMs [15, 16, 11, 17] to provide compact, relational representations that enable more interpretable reasoning [18, 19, 20]. However, existing approaches largely rely on static, corpus-wide graphs, which may not adapt well to the specific reasoning needs of individual queries. Constructing and maintaining large-scale KGs also incurs substantial costs in open-domain settings.
To address these challenges, we propose Retrieval-And-Structuring (RAS), a framework that dynamically constructs and reasons over query-specific knowledge graphs through iterative retrieval and structured knowledge building. RAS interleaves three key components: (1) a planning module that strategically identifies knowledge gaps and generates targeted sub-queries, (2) a retrieval-and-structuring module that extracts factual triples from retrieved passages and incrementally builds a query-specific graph, and (3) a knowledge-augmented answering module that generates outputs conditioned on the accumulated structured knowledge.
RAS addresses several critical limitations of prior methods. Unlike traditional RAG approaches that rely on single-pass retrieval [8, 9, 10], RAS employs iterative planning to actively identify and fill knowledge gaps during inference. In contrast to static KG-based approaches [11, 17], RAS dynamically constructs query-specific knowledge graphs tailored to the evolving reasoning needs of each query. This targeted construction captures only task-relevant information, avoiding the inefficiencies and noise associated with static, corpus-level graphs, which enables more precise and robust language generation.
@@
Through extensive evaluations across seven benchmarks spanning open-domain QA, closed-set QA, and long-form generation, RAS consistently outperforms strong baselines by 6.4% with open-source LLMs and 7.0% with proprietary models. Our main contributions are:
@@
We introduce RAS, a novel framework for dynamic knowledge structuring through iterative retrieval and graph construction, tailored to each query.
We develop a structure-aware planning and answering model that operates over evolving knowledge graphs.
@@
We demonstrate RAS’s effectiveness through comprehensive experiments and ablation studies across diverse knowledge-intensive tasks.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) [8, 9] enhances language model performance on knowledge-intensive tasks by incorporating retrieved passages into the model input, improving factual accuracy and grounding. Early RAG approaches retrieved a fixed number of passages once before generation [21, 22, 23], while later methods explored adaptive retrieval [24], where models dynamically issue retrieval queries based on uncertainty signals, or retrieval evaluation [25] to improve relevance. However, these techniques often struggle with complex or multi-hop reasoning due to their linear designs. Recent advances aim to strengthen retrieval quality by iterative refinement. Iterative retrieval-generation approaches [21, 26] leverage historical context to guide subsequent retrievals, while targeted retrieval strategies based on subqueries [27] or follow-up questions [28, 29] progressively enrich the retrieved evidence. Self-RAG [30] introduced self-reflective retrieval, allowing LLMs to selectively query for missing information and evaluate the quality of retrievals. RPG [31] further refined retrieval by extracting fine-grained paragraphs to enhance relevance. Despite these advances, existing methods face persistent challenges: retrieved context often contains redundant or irrelevant information, or conversely, discards essential facts during refinement. To address these issues, we propose converting retrieved content into a structured graph format and maintaining an iteratively enriched, query-specific knowledge graph, allowing the model to reason over an organized and progressively expanded knowledge structure.
Graph as Context for LLMs. Integrating structured graphs into language model workflows has emerged as a powerful approach for enhancing reasoning and factual accuracy [32, 33, 34, 35, 36, 37]. Graphs provide explicit, relational representations that help models go beyond flat text by making entity connections and multi-hop relationships more tractable. Recent work, such as GraphToken [38], shows that LLMs can process serialized graph inputs through token-level prompts [39] without modifying model parameters. G-Retriever [11] leverages global knowledge graphs (KGs) to retrieve entity-centric subgraphs, while GraphRAG-style methods [17, 40] construct corpus-level graphs and generate summaries to enhance retrieval quality. Despite their benefits, these methods rely heavily on static, global graphs, which often include irrelevant or redundant information and cannot adapt to the needs of individual queries. Moreover, building and maintaining large-scale KGs is computationally expensive, particularly in open-domain settings. This motivates a shift toward dynamic, inference-time graph construction. Our approach embraces this direction by proposing a retrieval-and-structuring framework that builds query-specific knowledge graphs on the fly. Rather than relying on fixed knowledge coverage, we iteratively retrieve and structure information into task-relevant graphs that evolve with the reasoning process. This dynamic graph construction aligns with recent interpretability findings [14], which suggest that LLM failures often stem from breakdowns in implicit reasoning chains—failures that explicit graph structuring can help mitigate.
3 Retrieval-And-Structuring (RAS) Framework
Effective knowledge-intensive language generation requires not only retrieving relevant information, but also structuring and reasoning over it systematically. We introduce Retrieval-And-Structuring (RAS), a framework that interleaves iterative retrieval planning with dynamic knowledge graph construction, enabling large language models (LLMs) to reason over progressively organized knowledge tailored to each query. Figure 1 illustrates the overall workflow.
Key Definitions. We define the core concepts used in the RAS framework as follows. The Main Query () denotes the original task input. A Subquery () is a focused retrieval query generated at iteration to obtain supporting evidence. Retrieved Passages () are the top- documents selected for each subquery. A text-to-triples model converts retrieved text into Extracted Triples (), structured as subject-predicate-object facts. These triples are incrementally accumulated into an evolving Knowledge Graph (), representing organized evidence about . The model produces an Action Plan () at each step, determining whether to continue retrieval ([SUBQ]), stop retrieval ([SUFFICIENT]), or, initially, answer directly without retrieval ([NO_RETRIEVAL]).
3.1 Knowledge State-Aware Planning
The Planning stage initiates and controls the retrieval-and-structuring process by dynamically assessing the current knowledge state.
@@ 3.1 Knowledge State-Aware Planning
Initial Planning. Formally, given an input query , the model generates an initial plan :
where is the planning instruction (as shown in Figure 15). can take one of two forms:
@@
[NO_RETRIEVAL]: If determines that can be answered directly without requiring any additional knowledge, the planning process terminates, and the framework proceeds directly to the final Answering stage (§3.3).
Iterative Planning. At iteration , given the accumulated knowledge and the subquery-triples history , the model updates the plan:
@@
where GNN is a graph neural network for encoding and projecting the evolving KG ; is the subquery at iteration , and is the extracted graph information (a list of triples) from the retrieved context .
The output at each iteration can be either:
@@
[SUBQ] : The model generates a new subquery to guide the retrieval of additional relevant knowledge. The subquery is designed to fill specific gaps in the current knowledge state with respect to answering . The framework proceeds to the next stage (§3.2).
[SUFFICIENT]: The accumulated knowledge is deemed sufficient to comprehensively address the main query . The iterative retrieval process terminates, and the framework proceeds to the Answering stage (§3.3).
@@
The Planning stage serves as a key driver of the RAS framework’s iterative retrieval and refinement process. By dynamically assessing the adequacy of the retrieved knowledge and generating targeted sub-queries, it enables the efficient acquisition of query-relevant information.
3.2 Text Retrieval and Knowledge Structuring
@@
This stage transforms retrieved text into structured knowledge and progressively builds the query-specific graph .
Text Retrieval. We use a text retriever to retrieve the top- semantically relevant passages from the corpus for each subquery :
We use a standard dense retriever by default but note that RAS is compatible with more advanced information retrieval methods [41, 42, 43].
Text-to-Triples (Graph) Conversion. To extract essential factual information from the retrieved passages , we employ a text-to-triples model . This model is trained on the full WikiOfGraph dataset [44], which is a high-quality, LLM-curated text-to-triples corpus. Details of the training process are provided in Appendix C.1. The model generates structured triples in the following format:
@@
where each triple represents a subject-predicate-object fact extracted from the text. This structured representation enables efficient downstream reasoning and facilitates integration with external knowledge graphs. Although is a lightweight LLM capable of fast inference using techniques such as quantization [45] and optimized inference frameworks like vLLM [46], the text-to-triples conversion can also be precomputed offline when maximal efficiency is required.
@@
Iterative Knowledge Enrichment to Query-Specific KG. The extracted triples are then converted into a graph structure , where and denote the sets of nodes and edges, respectively. Each node corresponds to a unique subject or object entity in , while each edge represents a predicate connecting two entities. To enrich the graph with semantic information, the attributes of nodes and edges are obtained through a text encoder (Sentence-BERT [47]):
@@
To progressively enrich the query-related knowledge in response to the evolving sub-queries, the structured graph at each iteration is merged into an evolving KG specific to the main query :
After enriching with the new knowledge, the Planning module (§3.1) reassesses the current knowledge to determine the next step. Based on and the chain of previous subqueries and their associated graph information, the model decides whether to generate another focused subquery for additional retrieval or to proceed with answering (§3.3) if the accumulated knowledge is sufficient.
3.3 Knowledge-Augmented Answering
At the final stage, the model generates an answer to the main query either conditioned on knowledge graph and subquery chain when retrieval-and-structuring was processed, or directly when no retrieval is needed.
@@ 3.3 Knowledge-Augmented Answering
where is the answering instruction (as shown in Figure 16). attends to knowledge in and subquery chain to generate accurate, coherent answers. This structured conditioning enables systematic reasoning grounded in the assembled knowledge.
3.4 Sturcture-Aware Multitask Learning
The RAS framework is trained through a multitask setup that unifies knowledge planning and knowledge-augmented answering under a standard next-token prediction objective.
Each training instance corresponds to either a planning task or an answering task, selected randomly:
Planning Task: Input: (encoded ), planning instruction , subquery-triples history, main query . Output: next plan .
Answering Task: Input: (encoded ), answering instruction , full subquery-triples history, main query . Output: final answer .
The model used in RAS is Graph LLM, an architecture adapted from prior work [38, 11]. It is trained to perform both planning and answering tasks via standard next-token prediction, conditioned on the instruction type and input structure. To the best of our knowledge, this is the first application of Graph LLMs to general language generation tasks such as open question answering.